Joes team can completely manicure the lawn at Detroit Park in 12 hours. When Sally's team was brought on to help, the job was done in about 5 1/2 hours. How long does it takes Sally's team to completely manicure the lawn when they are working alone?
Answer by stanbon(75887)

Joes team can completely manicure the lawn at Detroit Park in 12 hours. When Sally's team was brought on to help, the job was done in about 5 1/2 hours. How long does it takes Sally's team to completely manicure the lawn when they are working alone?
----
Joe's rate: 1/12 job/hr
Together rate: 1/(11/2) = 2/11 job/hr
Sally rate = 1/x job/hr
-----
Equation:
rate + rate = together rate
1/12 + 1/x = 2/11
---
Multiply thru by 132x to get:
11x + 132 = 24x
13x = 132
x = 10 hrs 9 minutes (time for Sally's team to do the job alone)
==========
Cheers,
Stan H
==========